Investors require a 17% rate of return on Brooks Sisters' stock (rs = 17%). What would the estimated value of Brooks' stock be if the previous dividend was D0 = $3 and if investors expect dividends to grow at a constant annual rate of (1) - 5%, (2) 0%, (3) 2%, or (4) 10%? Round your answers to the nearest cent.
Answer #1

Current price=D1/(Required return-Growth rate)

1.Current price=(3*(1-0.05))/(0.17-(0.05)

=2.85/0.22=$12.95(Approx).

2.Current price=3/0.17=$17.65(Approx)

3.Current price=(3*1.02)/(0.17-0.02)=$20.4

4.Current price=(3*1.1)/(0.17-0.1)=$47.14(Approx).